OpenStudy (anonymous):

The volume of a pyramid with a rectangular base is V=1/3LWH . Find W if V = 500, L = 15, and H = 4. A. W = 25 B. W = 20 C. W = 18 D. W = 15

OpenStudy (hoblos):

V=1/3LWH then W = 3V/LH = 3(500)/(15)(4) = 25